# Survey of Natural Language Processing for Education: Taxonomy, Systematic Review, and Future Trends

### v3

@@ 3.2.2 Distractor Generation
As we mentioned above, DG is another component for question construction, which aims to generate distractors with the given context, question, and correct answer. A good distractor should be relevant to the context, grammatically coherent to the interrogative of the question, and deceptive as a false answer [132].
@@
Datasets. The aforementioned MCQs for QG tasks such as SciQ and RACE can also be leveraged for DG tasks. Specifically, the distractors in SciQ are designed by crowd workers with reference [132]. Instead of providing entities and concepts as distractors, the distractors of RACE are at sentence-level [148], which are collected from the secondary school English exams. Alongside the datasets mentioned before, there are other MCQ datasets especially designed for DG. The data of MCQL [68] is crawled from the web and provides distractors that are either phrases or sentences, covering multi-subjects. Besides, Televic [149] is collected through Televic Education’s platform with distractors manually crafted by experts, featuring distractors across multiple languages and domains.
@@
Methods. Early studies for DG are rule-based and feature-based, which rely on linguistic rules [64] and specific features like word frequency [65] or diverse similarity measurement between distractors and correct answers [68]. With the rapid development of deep learning, the construction of distractors can be divided into two lines of mainstream: generation-based and ranking-based categories [70]. Generation-based methods automatically generate distractors token by token and ranking-based methods regard DG as a ranking task for a pre-defined distractor candidate set. Amongst the generation-based methods, Gao et al. [66] proposed a hierarchical model with dual attention for sentence and word relevance and irrelevant information filtering to generate distractors that are semantically consistent with and traceable in the context. Besides, Bitew et al. [67] employed LLMs for DG. They feed relevant questions as in-context examples to generate more plausible distractors. As for the ranking-based method, Liang et al. [68] proposed a two-stage framework: the first stage trains a model as a filter to reduce the candidate size, and the second stage trains another model for ranking the candidate. However, the two-stage framework encounters error accumulation, and some studies have discarded the first stage to address this issue, but led to reduced efficiency [69]. As follow-up studies, Wang et al. [70] proposed an end-to-end model for retrieval in the distractor candidates with a balance of efficiency and high performance.
@@ 3.3.1 Automated Essay Scoring
AES aims to develop a system to score an essay automatically, where the written text is given as the input and the system summarizes its quality as a score. The challenge of this task lies in understanding the content of the essay by modeling the different facets of the text. Until now, a series of methods have been developed to automatically score the essays.
Dataset Essay Types Source Language Exp #Q
CLC-FCE A/N/C/S/L non-native: ESOL test English holistic 1111K
ASAP A/R/N native: grades 7-10 exam English holistic 17171717K
TOEFL 11 A non-native: TOEFL test English holistic 1111K
ICLE A non-native: English organization/clarity/ 4444K
undergraduate exam adherence/persuasiveness
HSK A/N non-native: HSK test Chinese holistic 10101010K
TABLE III: Comparison of AES datasets. The essay types contain: argumentative (A), response (R), narrative (N), comment (C), suggestion (S) and letter (L). “Exp” denotes the type of scores annotated. “#Q” denotes the number of questions.
@@
Datasets. The Cambridge Learner Corpus-First Certificate in English exam (CLC-FCE) [151] is a widely used dataset for AES tasks. It provides both the holistic score and the manually tagged linguistic error types such that it makes possible to evaluate the instructional feedback of the developed AES systems. It is worth noting the dataset is annotated with the prompt of the essay, which indicates the topic of the essay. The Automated Student Assessment Prize (ASAP) is a Kaggle competition released for AES. A similar dataset TOEFL 11 corpus [152] contains essays from TOEFL exams, and it comes with a three-level proficiency annotation, which facilitates the AES research with the consideration of the difficulty level. Instead of holistic scores, fine-grained instructional annotation is needed for dimension-specific essay scoring research. International Corpus of Learner English (ICLE) and Argument Annotated Essays (AAE) [153] are two datasets containing essays annotated with the perspectives of structure organization, thesis clarity, prompt adherence, and argument persuasiveness. There are a few AES datasets, which are designed for other languages such as Chinese [154]. We display the details of the above datasets in Table III.
Methods. Due to the availability of the annotation and wide applicability of the commercial value, the majority of the studies usually predict a single score for an essay. The early AES systems can be formulated as a regression task, a classification task, or a ranking task. Specifically, a series of features such as length-based features [71], lexical features [72, 71], prompt-relevant features [73], readability features [71], syntactic features [71, 74] and so on, are first extracted from the essay, then off-the-shelf learning algorithms are leveraged to fit the annotated scores [155]. For regression, linear regression, support vector regression, and sequential minimal optimization are employed. For classification, SMO, logistic regression, and Bayesian networks are utilized. For ranking, SVM ranking can be applied. Subsequently, researchers propose neural networks with different architectures, including CNN, LSTM, and attention mechanism, to encode the essay and eventually predict the scores. Particularly, the document structure can be captured via hierarchically encoding the sentences from the word level to the sentence level. The supervision signals can be formulated with the combination of regression and ranking [75]. Hybrid methods are proposed to integrate the handcrafted features in deep learning-based methods [156]. Zeng et al. citezeng:aaai2023 applied curriculum learning to the AES training procedure.
Besides, there are some other challenges existing in the AES tasks. Due to the sparse annotation of the essays in the target domains, multiple studies [157, 158, 159] design neural networks decoupling the domain-dependent and domain-independent features, to facilitate the models to learn knowledge from different domains. Other methods aggregate multiple heuristic signals as supervision [79], enhance the features with capsule networks [78], and conduct interpretable essay scoring systems [77]. Regarding the evaluation of the essays, multiple dimensions are considered in a few studies such as cohesion, syntax, vocabulary, phraseology, grammar, and conventions[80, 81], which is able to construct more diverse evaluation metrics.
@@ 3.3.2 Automated Code Scoring
Similar to AES, ACS grades the score of a code snippet on various dimensions. Comparing with essay grading, code grading is more complicated in understanding the long-dependency inside a code snippet and the relationship between identifiers.
@@
Methods.
Code correctness is a vital factor in judging the quality of a code snippet, which can be quickly evaluated using compilers and interpreters. Therefore, dynamic analyses evaluate a code snippet by its performance of execution. By executing on a set of test cases, the execution results are supposed to be the same as the excepted results. In this way, many dynamic analysis tools for detecting code quality issues are developed, such as AutoStyle [82], WebTA [83], Refactor Tutor [84] and Hyperstyle [85]. On the contrary, some researchers apply static analysis approaches to assess a code snippet, where static features are extracted from the code without executing it. The static features include complexity, comments, number of lines, number of operators, number of operands and so on. The static analysis approaches leverage these features to identify the quality of code [86]. Machine learning algorithms also have been applied to ACS. Clustering, breadth-first search, and neural network approaches are utilized to track the logic errors in code [87, 88, 89]. Besides, Queiros et al. [90] proposed hybrid approaches combining static and dynamic approaches.
Instead of directly scoring a code snippet based on the extracted features, there are a series of studies focusing on evaluating the contextual embedding of a code snippet, which benefits to measure the distance between the evaluated code and the correct code for ACS [161]. The embedding of variable and method identifiers utilizes local and global context[162], abstract syntax trees [163], and AST paths [164, 165]. Later, Feng et al. [166] and Kanade et al. [167] proposed CodeBERT and CuBERT, respectively. They model code beyond the simple tokenization.
@@ 3.4.1 Grammatical Error Correction
GEC tasks are designed to identify and correct grammatical errors in text, including spelling errors caused by phonological confusion and visual confusion, as well as grammatical errors caused by incorrect use of grammatical rules. For the GEC task, how to correctly identify and correct grammar errors is difficult, and how to avoid over-correction is challenging.
@@
Datasets. Lang-8 is one of the most widely used datasets for language error correction, covering 80 languages, but its distribution of languages is very skewed, with Japanese and English being its most prevalent languages [168]. CLANG-8 [168] is a clean dataset based on Lang-8, fixing some mismatches between data sources and targets in Lang-8. CoNLL-2014 [169] and BEA-2019 [170] are both commonly used English language error correction datasets that refine different types of syntax errors. In addition, with the popularity of GEC tasks, more and more non-English datasets have been proposed and widely used by researchers. SIGHAN [171] is an early proposed data set for Chinese spelling check, which only contains grammatical errors such as pronunciation and shape errors of characters. The CTC [172] dataset not only contains spelling errors, but also introduces grammatical error and semantic error, which further improves the error types of the Chinese grammatical error dataset. The FCGEC [173] and FlaCGEC [174] datasets have added more linguistic annotations, introduced more granular grammatical rules, and covered 28 and 210 error types, respectively, presenting new challenges to the task of language error correction. At the same time, there are many GEC datasets for low-resource languages. There are dataset GECCC [175] developed in Czech, dataset RULEC-GEC [176] developed in Russian, dataset Falko-Merlin [177] developed in German, dataset COWS-L2H [178] developed in Spanish, dataset UA-GEC [179] developed in Ukrainian and dataset RONACC [180] developed in Romanian.
@@
Methods. Early work to handle language error correction tasks is mostly achieved by manually defining rules or building classification-based models, with modest performance [91, 92]. With the widespread use of neural networks, Seq2seq [181] model has become the mainstream method to solve the task. It treats grammatical error correction as a monolingual machine translation task, in which source sentences containing grammatical errors are tagged one by one to be translated into the correct target sentences. The Seq2seq methods are mainly built based on RNN, CNN, or Transformer framework. BART [93], T5 [94], EBGEC [95], SynGEC [96] and CSynGEC [97] are good models in practice. After that, Sequence-to-editing (Seq2edit)  [182] methods have gradually emerged. It treats the GEC task as a sequential labeling problem, marking the text spans with appropriate error tags, leaving the rest of the text unchanged, and generating one edit operation per prediction. The commonly used efficient Seq2edit models include PIE [98], EditNTS [99], LaserTagger [100], GECToR [101] and so on. Recently, the TemplateGEC [183] model proposed by Li et al. integrates the two frameworks of Seq2seq and Seq2edit, making use of their ability in error detection and correction, alleviating the problem of over-correction of seq2seq model to a certain extent, enhancing the effectiveness and robustness of the model. As a result, it gains impressive results in GEC tasks. Even though LLMs are able to solve GEC tasks in nature, researchers found that they encounter severe issues of over-correction [184].
Dataset Language Source Exp #Q
LANG-8 Multi LANG-8 Website 1M
(80 languages)
CLANG-8 En, GE, RU LANG-8 2.6M
CoNLL-2014 En NUCLE [185] 58K
BEA-2019 En NUCLE, FCE, W&I, 686K
LOCNESS, LANG-8
SIGHAN Zh - 12K
CTC Zh Texts written by native speakers 218K
FCGEC Zh Chinese examinations 41K
FlaCGEC Zh HSK [186] 13K
GECCC CS Essays, website texts 83K
RULEC-GEC RU - 12K
Falko-MERLIN GE MERLIN, FalkoEssayL2, 24K
FalkoEssayWhig
COWS-L2H ES Essays 12K
UA-GEC UK - 20K
RONACC RO TV and radio shows 10K
Defects4J Java Open source programs - 357
ManyBugs C Open source programs - 185
IntroClass C Student programs - 998
QuixBugs Java, Python Quixey Challenge  [^8] https://gigaom.com/2012/01/19/quixey-challenge/ - 40
Bugs2Fix Java Github - 2.3M
CodeReview Multi Github - 642K
(9 languages)
CodeReview-New Multi CodeReview, code reviews - 15K
(16 languages)
TABLE IV: Comparison of GEC and CEC datasets. “Exp” denotes whether the annotation contains the explicit error types. “#Q” denotes the number of sentences in the dataset.

### v4

@@ 3.2.2 Distractor Generation
As we mentioned above, DG is an associated component for question construction, which aims to generate distractors with the given context, question, and correct answer in a MCQ. A good distractor should be relevant to the context, grammatically coherent to the interrogative of the question, and deceptive as a false answer [151]. The task is generally formulated as follows:
$L_{\text{DG}}$ $=-\mathbb{E}_{(C,Q,A,O)\in\mathcal{D}}\log P_{\theta}(O\mid C,Q,A)$
Here, $C$, $Q$, AA, and $O$ are context, question, answer, and distractors, respectively.
@@
Datasets. The aforementioned MCQs for QG tasks such as SciQ and RACE can also be leveraged for DG tasks. Specifically, the distractors in SciQ are designed by crowd workers with reference [151]. Instead of providing entities and concepts as distractors, the distractors of RACE are at the sentence level [169], which are collected from the secondary school English exams. Alongside the datasets mentioned before, there are other MCQ datasets especially designed for DG. The data of MCQL [145] is crawled from the web and provides distractors that are either phrases or sentences, covering multi-subjects. Besides, Televic [146] is collected through Televic Education’s platform with distractors manually crafted by experts across multiple languages and domains.
@@
Methods. Early studies for DG are rule-based and feature-based, which rely on linguistic rules [170] and specific features like word frequency [171] or diverse similarity measurement between distractors and correct answers [145]. With the rapid development of deep learning, the construction of distractors can be divided into two lines of mainstream: generation-based and ranking-based categories [46]. Generation-based methods automatically generate distractors token by token and ranking-based methods regard DG as a ranking task for a pre-defined distractor candidate set. Hence, $O$ could be either generated or selected via generation-based or ranking-based models. Noting that there are some key challenges of DG to make a good distractor.
  • The produced distractors are irrelevant to the questions. Some studies [45, 46, 47] focus on generating distractors that are consistent with the context or the answer. Gao et al. [48] proposed a hierarchical model with dual attention for sentence and word relevance and irrelevant information filtering to generate distractors that are semantically consistent with and traceable in the context. Ding et al. [49] propose a multi-modal DG model integrating contrastive learning to ensure consistency and have high relevance to the contexts.
  • The distractors lack strong distractive effectiveness. A good distractor should be misleading [50], with high similarity to the correct answer and distinguishability from other distractors. Taslimipoor et al. [51] proposed a two-step approach using transfer learning to generate correct answers and distractors simultaneously, followed by clustering to classify and remove duplicates to avoid excessive similarity among distractors. Qu et al. [52] proposed an unsupervised DG framework using distillation from LLMs and contrastive decoding to avoid excessive similarity between generated distractors.
Our investigation reveals that DG, as a significant component of QG, has not been fully explored. More discussion on constructing DG datasets, improving evaluation strategies, and generating “good” distractors should be learned.
@@ 3.3 Automated Assessment
Dataset Essay Types Source Language Exp #Q Top 1 Top 2 Top 3
[…]
ESOL test Two-stream [172] Autom [172] Text-only [172]
[…]
grades 7-10 MTAA [55] Li&Ng [173] Chen&Li [58]
[…]
undergraduate exam /persuasiveness PMAES[174] Kumar[175] Uto[176]
[…]
grades 8128-12 MTAA [55] MTL-Vanilla [55] RoBERTa
[…]
HSK test DNN[178] LogReg[179] -
[…]
@@ 3.3.1 Automated Essay Scoring
AES aims to develop a system to score an essay automatically, where the written text is given as the input and the system summarizes its quality as a score:
$L_{\text{AES}}=-\mathbb{E}_{(E,s)\in\mathcal{D}}P_{\theta}(s|\mathbf{f}(E)),$ (1)
where $s$ is a holistic score and $\mathbf{f}(E)$ is a set of features derived from an essay $E$. The challenge of this task lies in understanding the content of the essay by modeling the different facets of the text.
@@
Datasets. The Cambridge Learner Corpus-First Certificate in English exam (CLC-FCE) [181] is a widely used dataset for AES tasks. Usually, a dataset is annotated with the prompt of the essay, which indicates the topic of the essay. Moreover, it provides both the holistic score and the manually tagged linguistic error types such that it makes possible to evaluate the instructional feedback of the developed AES systems. The Automated Student Assessment Prize (ASAP) is a Kaggle competition released for AES. A similar dataset ELLIPSE [182] contains essays from English language learners of the diverse difficulty levels. Instead of holistic scores, scores along different dimensions of essay quality are needed. International Corpus of Learner English (ICLE++) [174] is a dataset containing essays annotated with the perspectives of structure organization, thesis clarity, prompt adherence, and argument persuasiveness. There are a few AES datasets, which are designed for other languages such as Chinese [179]. We display the details of the above datasets in Table III.
Methods. The majority of the AES studies usually predict a single score for an essay. In Equation (1), $P_{\theta}(\cdot)$ can be formulated as a regression task, a classification task, a ranking task or their hybrid [183, 108]. A series of features such as length-based features [184], lexical features [185, 184], prompt-relevant features [186], readability features [184], syntactic features [184, 187] and so on, are first extracted from the essay, then off-the-shelf learning algorithms are leveraged to fit the annotated scores [188]. Besides the hand-crafted features, $\mathbf{f}(E)$ can be modules automatically capturing the document structure [176, 189]. In addition, there are some focuses existing in the AES tasks for education.
  • The singular evaluation of the essays lacks comprehensiveness. Multiple dimensions are considered in a few studies such as cohesion, syntax, vocabulary, phraseology, grammar, and conventions[53, 54], which are able to construct more diverse evaluation metrics. To enrich the evaluation of the AES task, researchers utilize LLMs to provide a more comprehensive and explainable evaluation of an essay [55]. With the trait-specific score, the AES system becomes multi-tasking learning:
    $L_{\text{AES}}=-\mathbb{E}_{(E,s)\in\mathcal{D}}\sum_{s^{k}\in s}P_{\theta}(s^{k}|\mathbf{f}(E)),$
    where $s^{k}$ measures different aspects of the essays.
  • The sparse annotation of the essays in domain-specific prompts. Cross-prompt AES aims to rate unseen target essays with annotated source prompts. Multiple studies [56, 57, 58] design neural networks decoupling the domain-dependent and domain-independent features, to facilitate the models to learn transitions ($s\rightarrow t$) from different domains: $t=s\times T^{\intercal}$, where $T$ is a matrix modeling the transition from source prompts to target prompts.
AES also faces the unsolved issues of object-oriented and scenarios-oriented scoring. Specifically, the essay scores of K12 learners, second-language learners and businessmen are different. Maybe a more fine-grained scoring paradigm or framework should be developed.
@@ 3.3.2 Automated Code Scoring
Similar to AES, ACS grades the score of a code snippet on various dimensions. Compared with essay grading, code grading is more complicated in understanding the long-dependency inside a code snippet and its execution effect could be a vital factor to judge.
@@
Methods. Code can be scored from different objectives, which is vital in judging the quality of a code snippet. Intuitively, by executing on a set of test cases, the execution results are supposed to be the same as the excepted results.
$\{\text{pass},\text{fail}\}=\text{Compiler}(C,\text{case}),$
where $C$ is a piece of code snippet. In this way, many dynamic analysis tools for detecting code quality issues are developed, such as AutoStyle [192], WebTA [193], Refactor Tutor [194] and Hyperstyle [190]. However, it is challenging to score a code snippet without executing it.
  • It is hard to identify good features and scoring algorithms. The static features include complexity, comments, number of lines, number of operators, number of operands, and so on. The static analysis approaches leverage these features to identify the quality of code [59]. Machine learning algorithms also have been applied to ACS. Clustering, breadth-first search, and neural network approaches are utilized to track the logic errors in code [60, 61, 62]. Besides, Queiros et al. [63] proposed hybrid approaches combining static and dynamic approaches, which can be formulated as follows:
    $L_{\text{ACS}}=-\mathbb{E}_{(C,s)\in\mathcal{D}}P_{\theta}(s|\mathbf{f}(C)),$
    where $C$ is the code snippet and $s$ is the annotated score for supervised training.
  • Implicitly encode the code snippet via distributed representation. Instead of directly scoring a code snippet based on the extracted features, there are a series of studies focusing on evaluating the contextual embedding of a code snippet, which benefits measuring the distance between the evaluated snippet and the correct snippet for ACS [64]. The embedding of variable and method identifiers utilizes local and global context[65], Abstract Syntax Trees (AST) [66], and AST paths [67, 68]. Later, Feng et al. [69] and Kanade et al. [70] proposed CodeBERT and CuBERT, respectively. They model the code beyond simple tokenization:
    $L_{\text{PreTrain}}=-\mathbb{E}_{C\in\mathcal{D}}P_{\theta}(C_{m}|C_{\backslash m}).$
    Similar to language models, the pre-training aims to learn the pattern of code language.
Even though a great number of systems have been developed for ACS, there still remains an unsolved question to develop adaptive and fine-grained scoring paradigms for ACS in various scenarios.
@@ 3.4 Error Correction
Dataset Language Source Exp #Q Top 1 Top 2 Top 3
LANG-8 Multi LANG-8 Website 11M - - -
(80 languages)
[…]
M-VOTING [73] GRECO [196] ESC [197]
[…]
LOCNESS, LANG-8 M-VOTING [73] GRECO [196] ESC [197]
[…]
native speakers SpellGCN [199] GECToR [200] SpellGCN [199]
[…]
LM-Combiner [72] STG [201] GPT2 [202]
[…]
EBGEC [205] BART [206] GECToR [200]
[…]
FalkoEssayWhig gT5 [207] BIFI [208] AG [209]
[…]
mT5 [210] Artificial [211] MT [212]
[…]
RoGPT2 [213] AT-GCM [214] NAT-GEC [215]
[…]
ThinkRepair [216] TENURE [217] KNOD [218]
[…]
SOSRepair [221] T5APR [219] Codex [222]
[…]
GenProg [224] TrpAutoRepair [225] FLEXIREPAIR [223]
[…]
ThinkRepair [216] BaseChatGPT [216] Codex [222]
[…]
AST-T5 [226] CodeT5 [227] PLBART [77]
[…]
(9 languages) LLaMA-Reviewer [228] CodeReviewer [229] CodeT5 [227]
[…]
(16 languages) code reviews ChatGPT CodeReviewer [229]
[…]
88footnotetext: https://gigaom.com/2012/01/19/quixey-challenge/
@@ 3.4.1 Grammatical Error Correction
GEC tasks are designed to identify and correct grammatical errors in text, including spelling errors caused by phonological confusion and visual confusion, as well as grammatical errors caused by incorrect use of grammatical rules.
$L_{\text{GEC}}=-\mathbb{E}_{(X,Y)\in\mathcal{D}}\log P_{\theta}(Y\mid X)$
Here $X$ and $Y$ are the source sentence and target sentence, respectively. For the GEC task, how to correctly identify and correct grammar errors is difficult, and how to avoid over-correction is challenging.
@@
Datasets. Lang-8 is one of the most widely used datasets for language error correction, covering 80 languages, but its distribution of languages is very skewed, with Japanese and English being its most prevalent languages [207]. CLANG-8 [207] is a clean dataset based on Lang-8, fixing some mismatches between data sources and targets in Lang-8. CoNLL-2014 [231] and BEA-2019 [232] are both commonly used English language error correction datasets that refine different types of syntax errors. In addition, with the popularity of GEC tasks, more and more non-English datasets have been proposed and widely used by researchers. SIGHAN [233] and CCTC [198] are Chinese GEC datasets introduced to support more CGEC research. The FCGEC [201] and FlaCGEC [204] datasets have added more linguistic annotations, introduced more granular grammatical rules, and covered $28$ and $210$ error types, respectively, presenting new challenges to the GEC task. At the same time, there are many GEC datasets for low-resource languages. There are datasets GECCC [209] in Czech, RULEC-GEC [234] in Russian, Falko-Merlin [235] in German, COWS-L2H [236] in Spanish, UA-GEC [237] in Ukrainian and RONACC [238] in Romanian.
@@
Methods. Early methods to handle GEC tasks are mostly achieved by manually defining rules or building classification-based models [239, 240]. With the widespread use of neural networks, Seq2seq [241] model has become the mainstream method to solve the task. It treats grammatical error correction as a monolingual machine translation task, in which source sentences containing grammatical errors are tagged one by one to be translated into the correct target sentences. The Seq2seq methods are mainly built based on RNN, CNN, or Transformer framework. BART [140], T5 [242], EBGEC [205], SynGEC [243] and CSynGEC [244] are good models in practice. After that, Seq2edit [245] methods have gradually emerged. It treats the GEC task as a sequential labeling problem, marking the text spans with appropriate error tags, leaving the rest of the text unchanged, and generating one edit operation per prediction. The commonly used efficient Seq2edit models include PIE [246], EditNTS [247], LaserTagger [248], GECToR [200] and so on. Recently, the TemplateGEC [249] model proposed by Li et al. integrates the two frameworks of Seq2seq and Seq2edit, making use of their ability in error detection and correction, alleviating the problem of over-correction of seq2seq model to a certain extent, enhancing the effectiveness and robustness of the model. While GEC methods have achieved high performance, it still faces some problems and challenges.
  • GEC with LLMs encounters over-correction issue. Even though LLMs are able to solve GEC tasks in nature, it has an over-correction issue, which hinders its application in the education domain since teachers are meant to correct the sentences with minimum edits. Some studies change the role of LLMs in GEC, utilizing them to improve the performance of the smaller model [71]. Also, some studies focus on addressing the over-correction issue in LLMs directly. Wang et al. [72] used cross-inference to generate a set of correction candidates and specifically train a model for soft ensemble within the candidate set to avoid over-correction. Omelianchuk et al. [73] used multiple models for voting to prevent over-correction in the output of a single model.
    $\hat{Y}=\text{Ensemble}_{Y_{i}\sim P_{\theta}(Y_{i}|X)}(Y_{i}).$
  • Distorted results of GEC systems in low resources. For language teaching over the world, the development of GEC systems for low-resource languages is underdeveloped. Many studies focus on applying current GEC methods to low-resource languages, such as Bangla [250], Slovak [251], Turkish [252], Zarma [253] and so on. Chan et al. [74] proposed a method for constructing multilingual data to handle mixed-language text, which significantly improves the performance of GEC systems.
Previous studies have not considered the GEC with visual features, where hand-written essays are submitted by students to be corrected. Besides, the evaluation criterion of GEC following minimum edits should be reconsidered in real-world since different corrections maybe acceptable.